# Conformal-in-the-Loop for Learning with Imbalanced Noisy Data

### v1

@@ Abstract
Class Imbalance Learning
Noisy Label Learning
Uncertainty Quantification
Conformal Prediction
journal: arXiv
@@ 1 Introduction
These interconnected challenges highlight the need for a more robust approach to training DNNs on imbalanced and noisy data. To address this, we introduce Conformal-in-the-Loop (CitL), a novel uncertainty-aware training method that dynamically adjusts sample weights and prunes data in response to uncertainty levels calculated through conformal prediction. Specifically, CitL applies conformal prediction to assess sample uncertainty, which is then used to strategically modify training in two primary ways: (1) Sample Weighting – assigning higher weights to samples with moderate uncertainty that are likely to be informative, thereby enhancing learning in minority classes, and (2) Sample Pruning – identifying and removing samples with high uncertainty that are likely mislabeled, reducing the noise that would otherwise degrade model performance. By leveraging these two mechanisms, CitL selectively refines the training data, ensuring the model focuses on reliable, informative samples. Through this targeted curation of data, CitL addresses the dual challenges of imbalance and noise while maintaining computational efficiency. Extensive experimental evaluations demonstrate that CitL significantly improves model performance across common data challenges of imabalanced, noisy labels. To summarize, our contributions in this paper are:
@@ 2.1 Learning with Class Imbalance
Class imbalance is a pervasive issue in many machine learning tasks, where the majority class is overrepresented in training data, often leading models to underperform on minority classes. This imbalance causes models to overpredict the majority class, resulting in frequent misclassification of minority instances, as noted in several studies [7, 6]. Solutions for addressing class imbalance generally fall into two main categories: data-centric and algorithmic approaches [29].
Data-centric methods focus on adjusting the training data distribution. Undersampling the majority class, though effective, risks discarding critical features, which can lead to a loss of valuable information. Clustering-based undersampling has been proposed to mitigate this by removing redundant or noisy samples while retaining informative examples [60, 34, 31, 5]. Oversampling the minority class can also help [51, 25], but it can increase the risk of overfitting and prolong training. Synthetic data generation techniques, such as SMOTE [36, 18], and data augmentation [61, 14] are commonly used to introduce new information, effectively balancing the data without merely duplicating samples.
@@ 2.2 Learning with Noisy Labels
A foundational work in the field of noisy label learning is Confident Learning (CL), a model-agnostic meta-learning algorithm designed to identify and remove noisy data before training. CL operates in two main phases: first, it estimates class-conditional label noise error rates using cross-validation; second, it employs a rank-and-prune strategy to eliminate the most problematic examples based on loss criteria from the cross-validated model [13]. CL’s approach includes data pruning through co-teaching, which assesses the generalizability of the dataset—if a dataset lacks generalizability, it indicates the presence of incorrect labels. The algorithm has also been applied iteratively iteratively [38]. After identifying and removing noisy samples, CL re-weights the loss for each class to account for the eliminated examples and re-trains on the complete dataset [37]. CL is predicated on the assumption that while the labels may be incorrect, the input data is valid, allowing for some useful information to remain as long as the class-conditional probability of noise, $\eta$, is less than or equal to 0.5. As $\eta$ approaches 0.5, training demands more data and time [3]. In essence, CL reverses the concept of boosting by downweighting or excluding high-loss examples, positing that these difficult instances may distort the final model.
@@ 2.3 Uncertainty Quantification
Uncertainty quantification is important in machine learning for understanding the variability in model predictions [20]. This variability comes from two main types of uncertainty: epistemic and aleatoric. Epistemic uncertainty arises from a model’s limited knowledge and can be reduced by adding more data or improving the model. In contrast, aleatoric uncertainty is the inherent randomness in the data that cannot be reduced [23]. Various methods have been developed to quantify these uncertainties, including Single Deterministic Networks, Bayesian Neural Networks (BNNs), and test-time augmentation techniques.
Single deterministic networks provide a clear framework for uncertainty estimation using a fixed architecture, allowing predictions through distributions like softmax in classification tasks. However, these networks often struggle with overconfidence in their predictions. To enhance uncertainty quantification, techniques such as Evidential Deep Learning [48] and Prior Networks [33] have been developed. These methods leverage higher-order distributions, with evidential learning predicting parameters of an evidential distribution and prior networks imposing priors on output distributions for better-calibrated uncertainty estimates. In contrast, BNNs address uncertainty by representing model parameters probabilistically, capturing uncertainties through posterior distributions. However, practical implementations often rely on approximations due to the computational intensity of exact Bayesian inference. Monte Carlo Dropout (MCD) [22] is one such approximation, using dropout during inference to sample from the predictive distribution. While averaging multiple predictions provides uncertainty estimates, this requires numerous forward passes, making it time-consuming for large networks. Additionally, selecting appropriate prior distributions for model parameters is critical, as incorrect choices can lead to miscalibrated uncertainty estimates. Similarly, test-time data augmentation generates multiple augmented samples to approximate predictive distributions, also incurring computational overhead. Thus, the effectiveness of these methods often involves a trade-off between uncertainty quantification quality and computational efficiency.
Recently, Conformal Prediction (ConP) has emerged as a robust framework for uncertainty quantification, offering confidence measures around predictions, which are particularly valuable for applications requiring reliable and interpretable outputs [2, 21]. Unlike traditional methods that often rely on assumptions about data distribution or specific model architectures, ConP is model-agnostic and distribution-free. This versatility allows it to be applied to any pre-trained model, whether it be a conventional machine learning algorithm or a complex deep neural network. Consequently, it is suitable for both classification and regression tasks across diverse fields.
@@ 2.4 Our Contribution
Fig. 1: An overview of the proposed Conformal-in-the-Loop (CitL) framework. Conformal Prediction is used to generate prediction sets during self-supervised machine learning. Training examples are weighted by uncertainty; validation examples are used to calibrate the uncertainty model.
@@ 3.1 Proposed CitL Method
Fitting the conformal classifier introduces memory and computational complexity proportional to the size of the calibration set, making it advantageous to keep this set as small as possible while still ensuring representative sampling of the data. During training, and once the conformal classifier is calibrated in the validation step, computational dependency shifts to the mini-batch size, which is typically manageable in most setups. This design allows efficient scaling, as the memory demands in subsequent steps remain relatively low.
@@ 3.2 Evaluation Benchmarks
For our evaluation benchmarks, we focus on two tasks: multiclass classification and semantic segmentation. For multiclass classification, we select CIFAR-10 [30] since it is known to be well-balanced; we then synthetically introduce label noise and class imbalance. For semantic segmentation, we select the CityScapes dataset [15], which is known to be imbalanced.
@@ 3.3 Experimental Setup and Implementation Details
Other Details. We implement our method, ConP, using the Least Ambiguous Set-Valued Classifier (LAC) [46] in PyTorch, leveraging GPU acceleration for enhanced computational efficiency. Data augmentation is performed using the Albumentations library [11], primarily aimed at preventing model overfitting. For CIFAR-10, we apply augmentations such as Horizontal Flip (50%), Crop and Pad (ranging from 0 to 56 pixels), and Coarse Dropout (covering a 56x56 area). For the CityScapes dataset, augmentation begins with a random scaling of the image size by up to ±10%, followed by a potential rotation of up to 10 degrees occurring 50% of the time. The image is then randomly cropped to a fixed size of 512x1024 pixels, with a 50% probability of horizontal flipping, and there is a 25% chance of applying one of several transformations, including reducing image quality to 90%, applying a blur, or adding Gaussian noise, along with brightness and contrast adjustments performed in 25% of cases.
@@ 3.4 Training Setup
The training environment includes a devcontainer with Python v3.12, PyTorch Lightning v2.4.0 [42], PyTorch v2.4 [4], and CUDA 12.1.0. Classification tasks are performed using MNasNet-Small [52] from the timm library [55], while the segmentation task employs EfficientNet-b0 [53] in a U-Net [45] configuration from Segmentation Models PyTorch [59]. Classification models are trained from scratch, whereas segmentation models use transfer learning from pretrained ImageNet weights. The best model to evaluate against the test set is selecting using minimum validation loss for baseline models. When using the method, the dynamic loss reweighting means that loss values cannot be effectively compared against epochs, so we select the best model by maximum validation accuracy for classification and mIoU for segmentation.
The learning rate is fixed at $5\times 10^{-4}$ and is reduced by a factor of 0.2 upon plateau, with a patience of 10 epochs. The minimum learning rate is set to $10^{-6}$, and weight decay is configured at 10% of the initial learning rate. Early stopping [40] is applied based on validation loss with a patience of 20 epochs. The ADAM optimizer [28] is used with default parameters $\beta=(0.9,0.999)$. All experiments are conducted on a single A6000 GPU with 256GB of system RAM.
@@ 4.1 Preliminary Experiment: Selecting the $\alpha$ Value
Fig. 2: Classification accuracy on imbalanced CIFAR-10 using our CitL method and the baseline with cross-entropy loss (blue dotted line) and focal loss (orange dotted line). The box plots show 10 $\alpha$ values ranging from 0.10 to 0.19.
Fig. 3: Segmentation mIoU for values of $\alpha$ [0.01, 0.10]. The best mIoU of 71.1 is achieved at $\alpha=0.02$. Baseline mIoU (66.1) is represented by the dotted blue line.
@@
Fig. 4: Segmentation mIoU on the validation set for various values of $\alpha$. The baseline performance is indicated by the dotted blue line.
@@ 4.3 Segmentation Results
Table 3: Classwise IoU for the CityScapes dataset with the best training run using $\alpha=0.02$. The table shows the baseline performance, performance with the method applied, and the resulting improvement for each class.
Baseline Method $\Delta$
Road 97.3 97.5 0.2
Sidewalk 80.3 81.3 1.0
Building 90.1 90.8 0.7
Wall 45.6 45.1 -0.5
Fence 47.5 51.3 3.8
Pole 58.1 60.9 2.8
Traffic light 57.9 64.3 6.3
Traffic sign 72.0 74.3 2.3
Vegetation 91.3 91.8 0.5
Terrain 56.9 58.5 1.6
Sky 93.2 94.1 0.9
Person 76.7 78.8 2.2
Rider 50.9 55.0 4.1
Car 92.8 93.5 0.7
Truck 49.0 66.7 17.7
Bus 48.0 66.0 18.1
Train 39.6 59.6 20.1
Motorcycle 36.7 47.8 11.1
Bicycle 71.9 73.8 1.9
Overall 66.1 71.1 5.0
@@ 4.4 Computational Overhead
Fig. 5: Average per training-step time of CitL over baseline. Multiple CitL values represent different $\alpha$ hyperparameters. (CIFAR-10 11.2 $\pm$ 0.8%, CityScapes 4.0 $\pm$ 0.4%)
@@
Fig. 6: Weight Range, showing the difference in mean weights between the highest-weighted and lowest-weighted classes on the CIFAR-10 dataset, stratified by noise level. Only the best performing $\alpha$ value is shown for each noise level.
Fig. 5 and Table 4 present an analysis of the overhead associated with our method across various tasks. Notably, classification tasks demonstrate a higher overhead compared to segmentation tasks, likely because baseline training steps are faster. However, the overall increase in overhead remains small, with a mean increase of 11.2% in training time per step for classification and 4.0% for segmentation. These findings are encouraging and suggest that our method is likely to scale effectively to larger tasks and datasets.
@@ 4.5 Example Weighting
This section presents an analysis of the weighting generated by our method on the imbalanced CIFAR-10 dataset and the CityScapes dataset. The primary objective is to demonstrate the method’s effectiveness in addressing dataset imbalances by emphasizing challenging examples.
In classification tasks, Fig. 6 illustrates the variation in mean weight between the highest- and lowest-weighted classes across different training steps. Notably, weighting is highest during the early training epochs and decreases as training progresses, reflecting the model’s evolving uncertainty, as shown in Fig. 8. In the absence of noise, the model successfully reduces uncertainty to zero; however, this complete reduction is unattainable in the presence of noise. Similarly, noise inhibits the method’s ability to diminish the weighting effect to zero.
@@
Fig. 7: Mean weight by class across training epochs for the imbalanced CIFAR-10 dataset with no synthetic label noise. Each subfigure shows the distribution of mean class weights at different training epochs, illustrating how the algorithm adjusts class weights over time.
Fig. 8: Uncertainty by noise level. Uncertainty is defined as the percentage of examples whose prediction set contains more than one possible class. With no noise, the model is able to reduce uncertainty to 0; noise makes this reduction to 0 impossible.
@@
Fig. 9: Mean weight by class across training epochs for the CityScapes dataset for best value of $\alpha$ = 0.02. Each subfigure shows the distribution of mean class weights at different training epochs, illustrating how the algorithm adjusts class weights over time.

### v2

@@ Abstract
Class Imbalance Learning , Noisy Label Learning , Uncertainty Quantification , Conformal Prediction ,
journal: Pattern Recognition
@@
[1]organization=University of British Columbia, addressline=, city=Kelowna, postcode=, state=British Columbia, country=Canada
[…]
[2]organization=University of Victoria, addressline=, city=Victoria, postcode=, state=British Columbia, country=Canada
[…]
[3]organization=University of British Columbia, addressline=, city=Vancouver, postcode=, state=British Columbia, country=Canada
@@ 1 Introduction
These interconnected challenges highlight the need for a more robust approach to training DNNs on imbalanced and noisy data. To address this, we introduce Conformal-in-the-Loop (CitL), a novel uncertainty-aware training method that dynamically adjusts sample weights and prunes data in response to uncertainty levels calculated through conformal prediction [25]. Specifically, CitL applies conformal prediction to assess sample uncertainty, which is then used to strategically modify training in two primary ways: (1) Sample Weighting – assigning higher weights to samples with moderate uncertainty that are likely to be informative, thereby enhancing learning in minority classes, and (2) Sample Pruning – identifying and removing samples with high uncertainty that are likely mislabeled, reducing the noise that would otherwise degrade model performance. By leveraging these two mechanisms, CitL selectively refines the training data, ensuring the model focuses on reliable, informative samples. Through this targeted curation of data, CitL addresses the dual challenges of imbalance and noise while maintaining computational efficiency. Extensive experimental evaluations demonstrate that CitL significantly improves model performance across common data challenges of imabalanced, noisy labels. To summarize, our contributions in this paper are:
@@ 2.1 Learning with Class Imbalance
Class imbalance is a pervasive issue in many machine learning tasks, where the majority class is overrepresented in training data, often leading models to underperform on minority classes. This imbalance causes models to overpredict the majority class, resulting in frequent misclassification of minority instances, as noted in several studies [3, 2]. Solutions for addressing class imbalance generally fall into two main categories: data-centric and algorithmic approaches.
Data-centric methods focus on adjusting the training data distribution. Undersampling the majority class, though effective, risks discarding critical features, which can lead to a loss of valuable information. Clustering-based undersampling has been proposed to mitigate this by removing redundant or noisy samples while retaining informative examples [39]. Oversampling the minority class can also help [34, 19], but it can increase the risk of overfitting and prolong training. Synthetic data generation techniques, such as SMOTE [13], and data augmentation [40, 9] are commonly used to introduce new information, effectively balancing the data without merely duplicating samples.
@@ 2.2 Learning with Noisy Labels
A foundational work in the field of noisy label learning is Confident Learning (CL), a model-agnostic meta-learning algorithm designed to identify and remove noisy data before training. CL operates in two main phases: first, it estimates class-conditional label noise error rates using cross-validation; second, it employs a rank-and-prune strategy to eliminate the most problematic examples based on loss criteria from the cross-validated model [8]. CL’s approach includes data pruning through co-teaching, which assesses the generalizability of the dataset—if a dataset lacks generalizability, it indicates the presence of incorrect labels. The algorithm has also been applied iteratively iteratively [24]. After identifying and removing noisy samples, CL re-weights the loss for each class to account for the eliminated examples and re-trains on the complete dataset [24]. CL is predicated on the assumption that while the labels may be incorrect, the input data is valid, allowing for some useful information to remain as long as the class-conditional probability of noise, $\eta$, is less than or equal to 0.5. As $\eta$ approaches 0.5, training demands more data and time. In essence, CL reverses the concept of boosting by downweighting or excluding high-loss examples, positing that these difficult instances may distort the final model.
@@ 2.3 Uncertainty Quantification
Uncertainty quantification is important in machine learning for understanding the variability in model predictions [15]. This variability comes from two main types of uncertainty: epistemic and aleatoric. Epistemic uncertainty arises from a model’s limited knowledge and can be reduced by adding more data or improving the model. In contrast, aleatoric uncertainty is the inherent randomness in the data that cannot be reduced. Various methods have been developed to quantify these uncertainties, including Single Deterministic Networks, Bayesian Neural Networks (BNNs), and test-time augmentation techniques.
Single deterministic networks provide a clear framework for uncertainty estimation using a fixed architecture, allowing predictions through distributions like softmax in classification tasks. However, these networks often struggle with overconfidence in their predictions. Techniques such as Evidential Deep Learning and Prior Networks have been developed to enhance uncertainty quantification [31]. These methods leverage higher-order distributions, with evidential learning predicting parameters of an evidential distribution and prior networks imposing priors on output distributions for better-calibrated uncertainty estimates. In contrast, BNNs address uncertainty by representing model parameters probabilistically, capturing uncertainties through posterior distributions. However, practical implementations often rely on approximations due to the computational intensity of exact Bayesian inference. Monte Carlo Dropout (MCD) [17] is one such approximation, using dropout during inference to sample from the predictive distribution. While averaging multiple predictions provides uncertainty estimates, this requires numerous forward passes, making it time-consuming for large networks. Additionally, selecting appropriate prior distributions for model parameters is critical, as incorrect choices can lead to miscalibrated uncertainty estimates. Similarly, test-time data augmentation generates multiple augmented samples to approximate predictive distributions, also incurring computational overhead. Thus, the effectiveness of these methods often involves a trade-off between uncertainty quantification quality and computational efficiency.
Recently, Conformal Prediction (ConP) has emerged as a robust framework for uncertainty quantification, offering confidence measures around predictions, which are particularly valuable for applications requiring reliable and interpretable outputs. Unlike traditional methods that often rely on assumptions about data distribution or specific model architectures, ConP is model-agnostic and distribution-free. This versatility allows it to be applied to any pre-trained model, whether it be a conventional machine learning algorithm or a complex deep neural network. Consequently, it is suitable for both classification and regression tasks across diverse fields  [16].
@@ 2.4 Our Contribution
Figure 1: An overview of the proposed Conformal-in-the-Loop (CitL) framework. Conformal Prediction is used to generate prediction sets during self-supervised machine learning. Training examples are weighted by uncertainty; validation examples are used to calibrate the uncertainty model.
@@ 3.1 Proposed CitL Method
Fitting the conformal classifier introduces memory and computational complexity proportional to the size of the calibration set, making it advantageous to keep this set as small as possible while still ensuring representative sampling of the data. During training, and once the conformal classifier is calibrated in the validation step, computational dependency shifts to the mini-batch size, which is typically manageable in most setups.
@@ 3.2 Evaluation Benchmarks
For our evaluation benchmarks, we focus on two tasks: multiclass classification and semantic segmentation. For multiclass classification, we select CIFAR-10since it is known to be well-balanced; we then synthetically introduce label noise and class imbalance. For semantic segmentation, we select the CityScapes dataset, which is known to be imbalanced.
@@ 3.3 Experimental Setup and Implementation Details
Other Details. We implement our method, ConP, using the Least Ambiguous Set-Valued Classifier (LAC) [30] in PyTorch, leveraging GPU acceleration for enhanced computational efficiency. Data augmentation is performed using the Albumentations library, primarily aimed at preventing model overfitting. For CIFAR-10, we apply augmentations such as Horizontal Flip (50%), Crop and Pad (ranging from 0 to 56 pixels), and Coarse Dropout (covering a 56x56 area). For the CityScapes dataset, augmentation begins with a random scaling of the image size by up to ±10%, followed by a potential rotation of up to 10 degrees occurring 50% of the time. The image is then randomly cropped to a fixed size of 512x1024 pixels, with a 50% probability of horizontal flipping, and there is a 25% chance of applying one of several transformations, including reducing image quality to 90%, applying a blur, or adding Gaussian noise, along with brightness and contrast adjustments performed in 25% of cases.
@@ 3.4 Training Setup
The training environment includes a devcontainer with Python v3.12, PyTorch Lightning v2.4.0, PyTorch v2.4, and CUDA 12.1.0. Classification tasks are performed using MNasNet-Small from the timm library, while the segmentation task employs EfficientNet-b0 in a U-Net configuration from Segmentation Models PyTorch. Classification models are trained from scratch, whereas segmentation models use transfer learning from pretrained ImageNet weights. The best model to evaluate against the test set is selecting using minimum validation loss for baseline models. When using the method, the dynamic loss reweighting means that loss values cannot be effectively compared against epochs, so we select the best model by maximum validation accuracy for classification and mIoU for segmentation.
The learning rate is fixed at $5\times 10^{-4}$ and is reduced by a factor of 0.2 upon plateau, with a patience of 10 epochs. The minimum learning rate is set to $10^{-6}$, and weight decay is configured at 10% of the initial learning rate. Early stopping is applied based on validation loss with a patience of 20 epochs. The ADAM optimizer is used with default parameters $\beta=(0.9,0.999)$. All experiments are conducted on a single A6000 GPU with 256GB of system RAM.
@@ 4 Results and Discussion
Figure 2: Segmentation mIoU on the validation set for various values of $\alpha$. The baseline performance is indicated by the dotted blue line.
Figure 3: Segmentation mIoU for values of $\alpha$ [0.01, 0.10]. The best mIoU of 71.1 is achieved at $\alpha=0.02$. Baseline mIoU (66.1) is represented by the dotted blue line.
@@ 4.2 Classification Results
Figure 4: Classification accuracy on imbalanced CIFAR-10 using our CitL method and the baseline with cross-entropy loss (blue dotted line) and focal loss (orange dotted line). The box plots show 10 $\alpha$ values ranging from 0.10 to 0.19.
@@ 4.3 Segmentation Results
Table 3: Classwise IoU for the CityScapes dataset with the best training run using $\alpha=0.02$.
Baseline Method $\Delta$ Road 97.3 97.5 0.2 Sidewalk 80.3 81.3 1.0 Building 90.1 90.8 0.7 Wall 45.6 45.1 -0.5 Fence 47.5 51.3 3.8 Pole 58.1 60.9 2.8 Traffic light 57.9 64.3 6.3 Traffic sign 72.0 74.3 2.3 Vegetation 91.3 91.8 0.5 Terrain 56.9 58.5 1.6 Sky 93.2 94.1 0.9 Person 76.7 78.8 2.2 Rider 50.9 55.0 4.1 Car 92.8 93.5 0.7 Truck 49.0 66.7 17.7 Bus 48.0 66.0 18.1 Train 39.6 59.6 20.1 Motorcycle 36.7 47.8 11.1 Bicycle 71.9 73.8 1.9 Overall 66.1 71.1 5.0
@@ 4.4 Computational Overhead
Figure 5: Average per training-step time of CitL over baseline. Multiple CitL values represent different $\alpha$ hyperparameters. (CIFAR-10 11.2 $\pm$ 0.8%, CityScapes 4.0 $\pm$ 0.4%)
@@
Fig. 5 and Table 4 present an analysis of the overhead associated with our method across various tasks. Notably, classification tasks demonstrate a higher overhead compared to segmentation tasks, likely because baseline training steps are faster. However, the overall increase in overhead remains small, with a mean increase of 11.2% in training time per step for classification and 4.0% for segmentation. These findings are encouraging and suggest that our method is likely to scale effectively to larger tasks and datasets. This section presents an analysis of the weighting generated by our method on the imbalanced CIFAR-10 dataset and the CityScapes dataset. The primary objective is to demonstrate the method’s effectiveness in addressing dataset imbalances by emphasizing challenging examples.
@@ 4.5 Example Weighting
For the classification task, Fig. 6(a) illustrates the variation in mean weight between the highest- and lowest-weighted classes across different training steps. Notably, weighting is highest during the early training epochs and decreases as training progresses, reflecting the model’s evolving uncertainty, as shown in Fig. 6(b). In the absence of noise, the model successfully reduces uncertainty to zero; however, this complete reduction is unattainable in the presence of noise. Similarly, noise inhibits the method’s ability to diminish the weighting effect to zero.
(a) Weight Range, showing the difference in mean weights between the highest-weighted and lowest-weighted classes on the imbalanced CIFAR-10 dataset, stratified by noise level. Only the best performing $\alpha$ value is shown for each noise level.
(b) Uncertainty by noise level for the imbalanced CIFAR-10 dataset. Uncertainty is defined as the percentage of examples whose prediction set contains more than one possible class. With no noise, the model is able to reduce uncertainty to 0; noise makes this reduction to 0 impossible.
@@
Figure 7: Mean weight by class across training epochs for the imbalanced CIFAR-10 dataset with no synthetic label noise. Each subfigure shows the distribution of mean class weights at different training epochs, illustrating how the algorithm adjusts class weights over time.
@@
Figure 8: Mean weight by class across training epochs for the CityScapes dataset for best value of $\alpha$ = 0.02. Each subfigure shows the distribution of mean class weights at different training epochs, illustrating how the algorithm adjusts class weights over time.